Exercise 12-5 Coburn (beginning capital, $60,000) and Webb (beginning capital $90,000) are partners. During 2017, the partnership earned net income of $80,000, and Coburn made drawings of $18,000 while Webb made drawings of $24,000. Assume the partnership income-sharing agreement calls for income to be divided 45% to Coburn and 55% to Webb, prepare the journal entry to record the allocation of net income. (Credit account titles are automatically indented when amount is entered. Do not indent manually.) Account Titles and Explanation Debit Credit Assume the partnership income-sharing agreement calls for income to be divided with a salary of $30,000 to Coburn and $25,000 to Webb, with the remainder divided 45% to Coburn and 55% to Webb, prepare the journal entry to record the allocation of net income. (Credit account titles are automatically indented when amount is entered. Do not indent manually.) Account Titles and Explanation Debit Credit

Assume the partnership income-sharing agreement calls for income to be divided with a salary of $40,000 to Coburn and $35,000 to Webb, interest of 10% on beginning capital, and the remainder divided 50%-50%. Prepare the journal entry to record the allocation of net income. (Credit account titles are automatically indented when amount is entered. Do not indent manually.) Account Titles and Explanation Debit Credit LINK TO TEXT Compute the partners ending capital balances under the assumption in part (c) above. Ending capital Coburn Webb
Answer #1

a) Profit and loss account . Dr. 80000

To Coburn Capital Account . (80000*45%) 36000

To Webb Capital account(80000*55%). 44000

(Being net income shared by 45% to coburn and 55% for webb)

b)Profit and loss appropriation account . Dr. 55000

To Coburn Capital Account . 30000

To Webb Capital Account. 25000

(Being salary allocated to parents out of profit)

Profit and loss appropriation account Dr . 25000

To Coburn Capital account.(25000*45%) 11250

To Webb Capital account (25000*55%). 13750

(Being remaining profit disturbed at 45% to coburn and 55% to Webb)

c) Profit and loss appropriation account. Dr. 75000

To Coburn Capital account. 40000

To Webb Capital account. 35000

(Being salary allocated to partners)

Profit and loss appropriation account . Dr . 5000

To Coburn Capital Account. (5000*6/15) . 2000   

To Webb Capital Account (5000*9/15). 3000

(Being interest in capital allocated in ratio of capital at the beginning because of insufficient funds)

Ending capital under assumption c) of:

  1. Coburn=60000-18000+40000+2000= $84000
  2. Webb=90000-24000+35000+3000=$104000